# Benchmarking Large Language Models for Cyberbullying Detection in Real-World YouTube Comments

### v1

@@ Abstract
As online platforms grow, comment sections increasingly host harassment that undermines user experience and well-being. This study benchmarks three leading large language models, OpenAI GPT-4.1, Google Gemini 1.5 Pro, and Anthropic Claude 3 Opus, on a corpus of 5 080 YouTube comments sampled from high-abuse threads in gaming, lifestyle, food vlog, and music channels. The dataset comprises 1 334 harmful and 3 746 non-harmful messages in English, Arabic, and Indonesian, annotated independently by two reviewers with substantial agreement (Cohen’s $\kappa$ = 0.83). Using a unified prompt and deterministic settings, GPT-4.1 achieved the best overall balance with an F1 score of 0.863, precision of 0.887, and recall of 0.841. Gemini flagged the highest share of harmful posts (recall = 0.875) but its precision fell to 0.767 due to frequent false positives. Claude delivered the highest precision at 0.920 and the lowest false-positive rate of 0.022, yet its recall dropped to 0.720. Qualitative analysis showed that all three models struggle with sarcasm, coded insults, and mixed-language slang. These results underscore the need for moderation pipelines that combine complementary models, incorporate conversational context, and fine-tune for under-represented languages and implicit abuse. A de-identified version of the dataset and full prompts is publicly released to promote reproducibility and further progress in automated content moderation.
@@
Artificial intelligence, content moderation, cyberbullying, hate speech detection, large language models, natural language processing
@@ I Introduction
Online platforms now mediate much of everyday communication, with more than 4.95 billion active social-media accounts worldwide in 2023 [1]. That ubiquity carries risk: recent studies report that between 6.5% and 35.4% of adolescents in the U.S. and Europe have experienced cyberbullying [2], and longitudinal studies link such exposure to elevated anxiety, depression, and suicidal ideation [3]. As the scale and psychological impact of online abuse grow, so does the urgency of developing reliable, context-aware moderation methods, especially for platforms popular with adolescents such as YouTube and TikTok.
@@ I-A Background
Cyberbullying has become a major digital-safety concern because abusive material is persistent, anonymous, and can spread instantaneously across networks [4]. As global connectivity and social-media adoption have expanded, online harassment has become more visible, with increasing concern around its scale and severity [5].
Comment sections on social networks, video-sharing sites, and forums are particular hotspots. In a 2024 U.S. survey of parents, 79 % said their children had encountered cyberbullying on YouTube [6]. A cross-national analysis of 180 000 adolescents in 42 countries also linked problematic social-media use to both cyberbullying victimisation and perpetration, with stronger effects among girls [7].
@@ I-B Motivation and Problem Statement
This is especially dangerous in cyberbullying, where the impact of a single overlooked comment can be deeply personal [12]. Unlike hate speech or spam, bullying is often subtle: expressed through mockery, group pressure, or repeated jabs that may seem harmless in isolation but accumulate over time. Many academic benchmarks fail to capture this complexity, relying on synthetic or crowd-sourced datasets with clear-cut abuse [13].
This study addresses that gap by evaluating three state-of-the-art LLMs, GPT-4, Gemini, and Claude, on authentic YouTube comments from videos with documented cyberbullying. By focusing on nuanced, emotionally loaded content, we examine not just whether models detect harmful language, but how well they handle ambiguity, cultural context, and tone.
@@ I-C Objectives of the Study
  1. 1.
    Benchmark performance. Compare OpenAI GPT, Google Gemini, and Anthropic Claude in detecting cyberbullying within real YouTube data.
  2. 2.
    Quantify error patterns. Measure false-positive and false-negative rates and identify recurring misclassification themes.
  3. 3.
    Conduct qualitative error analysis. Manually inspect model outputs to assess contextual understanding and pinpoint linguistic edge cases.
  4. 4.
    Assess robustness across nuance. Evaluate how well each model handles abusive content that is linguistically ambiguous, culturally specific, or indirectly expressed.
@@ II-A AI Content Moderation and Toxic Language Detection
However, most benchmarks are composed of isolated sentences and short comments labeled through crowdwork. They rarely reflect the conversational, multi-turn, or emotionally charged dynamics seen in real-world platforms. Furthermore, these datasets often overrepresent overt toxicity while underrepresenting gray-area content such as mockery, dismissive sarcasm, or personal digs. As a result, models trained on them perform well in benchmark settings but struggle in live deployments where language is more ambiguous and emotionally expressive [17, 18].
@@ II-B Challenges in Detecting Implicit and Nuanced Abuse
Detecting implicit harm remains one of the most difficult challenges in AI-based moderation. Sarcasm, coded language, cultural slang, and subtextual insults frequently evade model detection, even in large LLMs. For example, a phrase like “She’s always such a queen, right?” may be an insult depending on tone and context, but appears harmless in isolation. Work on sarcasm detection [19], adversarial misspellings, irony detection, and tone modeling has attempted to address this gap, but progress remains limited [20].
Recent research also suggests that models often overflag emotionally intense or sensitive language even when used in supportive contexts, leading to user frustration and moderation fatigue. This tension between underflagging subtle harm and overflagging emotional expression makes fine-tuned moderation especially difficult. Studies have proposed integrating tone analysis, conversational history [22], and speaker intent as ways to improve subtle abuse detection, but these are rarely deployed at scale.
II-C Moderating Multilingual and Underresourced Content
While English dominates most training corpora, abuse occurs in every language, and much of it is multilingual, transliterated, or mixed with emojis and slang. Studies show a consistent drop in model accuracy when applied to languages such as Arabic [23], Serbian, Tagalog, or Swahili, particularly when users apply spelling variation or obfuscation to evade filters. Code-switching and transliteration further degrade detection performance, as the model may miss common patterns or interpret them out of context [24, 25].
Cross-lingual models and prompt-tuned systems have shown moderate gains, especially when fine-tuned on translated or augmented datasets. However, language imbalance persists, especially in online spaces with small but highly active non-English user communities. For platforms with global audiences, this introduces a serious equity issue: harmful content in English is far more likely to be caught than equivalent abuse in low-resource languages or regional dialects.
@@ II-D Evaluating LLMs for Safety and Fairness in Moderation
The rise of general-purpose LLMs like GPT-4, Claude, and Gemini has led to growing interest in using them for content moderation. Some recent work shows that these models can outperform task-specific classifiers in zero-shot settings, especially on unseen data [13, 21]. They are also more flexible, capable of incorporating platform policies or contextual reasoning in their prompts. However, their moderation performance can be highly sensitive to prompt phrasing, target domain, and cultural framing.
More recent research has introduced sociocultural audit frameworks, including persona-based prompts [27] and synthetic demographic simulations, to evaluate fairness and blind spots. These studies highlight that model behavior often shifts depending on perceived user identity or topic domain, raising concerns about bias and overgeneralization [28]. While LLMs offer clear advantages in flexibility and scalability, their reliance on pretraining data and lack of transparency in decision-making still pose barriers to safe deployment in moderation workflows.
This study builds on these gaps by evaluating leading LLMs in a real-world moderation scenario: cyberbullying detection in emotionally charged YouTube comment sections, covering multiple languages and informal speech styles. It extends previous work by focusing not only on overall accuracy but on model behavior across sarcasm, cultural cues, and under-resourced content.
@@ III-B Data Cleaning and Anonymization
Each comment underwent a two–step preprocessing pipeline. First, text was normalised (UTF-8 decoding, whitespace trimming, emoji and punctuation were preserved). Second, personally identifiable information including user names, real names, phone numbers, e-mail addresses, links, and explicit geo-markers was either deleted or replaced by neutral placeholders such as UserNameProtected. No other linguistic content was altered. This procedure leaves the semantic core of each message intact while mitigating re-identification risk, allowing the corpus to be shared for research without exposing private data.
@@ III-C Data Labeling and Ground Truth
Inter‐rater reliability was high, with 91.0% raw agreement and $\kappa=0.83$, which is considered substantial. [^1] Landis and Koch, 1977. Disagreements (9 % of instances) were adjudicated by discussion until consensus, and that label was written to the final ground-truth file. Labels were never exposed to the models.
@@ III-G Manual Review and Model Comparison
Quantitative scores alone do not reveal why a system succeeds or fails, so we carried out a post-hoc qualitative review. Two annotators independently inspected a stratified sample of 200 disagreements, consisting of 50 false positives, 50 false negatives, and 100 edge-case ties, drawn in equal proportion from all three models. For each comment we discussed (i) whether the human ground truth should stand and (ii) what linguistic cues might have misled the model (sarcasm, coded slurs, emoji, topic drift, and so forth). Notes from this exercise were grouped into recurring error themes that inform the Discussion section. The procedure does not alter any numeric results but clarifies how each system handles nuance, context, and cultural references.
@@ IV-A Dataset Composition
The evaluation dataset consisted of 5 080 comments, including 1 334 harmful and 3 746 non-harmful instances. Harmful content made up about 26.3 percent of the total dataset, indicating a moderate class imbalance that could influence model metrics like precision and recall. Table I provides the class breakdown.
@@ V-A False Positives and Surface-Level Flagging
To better understand model overflagging, we analyzed cases where AI models marked comments as harmful while human reviewers did not. These false positives were grouped by model agreement patterns to highlight shared tendencies and specific weaknesses. While most flagged content reflected strong language or emotional tone, it often lacked any abusive intent. Below are four common patterns observed.
V-A1 All Three Models vs. Human: Shared Overflagging Patterns
@@ V-A3 Gemini: Over-Flagging Concern and Instability
Gemini frequently flagged comments that referenced mental health, personal change, or emotional concern, even when human reviewers found no clear harm. Many of these comments offered support, expressed worry, or questioned the authenticity of a user’s behavior in dramatic or sarcastic terms. Phrases like “please take care of yourself” or “you can’t keep scaring everyone like this” were often interpreted by Gemini as harmful. This suggests a tendency to overflag emotionally sensitive or speculative language, particularly when it touches on perceived instability or trauma.
@@ V-C Risks of Missed Moderation in Sensitive Cases
More troubling were coordinated suggestions to ”unfollow so he falls off” or ”ignore him until he cracks again.” These comments reflect collective hostility rather than isolated aggression. When such messages go unflagged, they enable mob harassment campaigns that can lead to reputational harm, social withdrawal, or emotional distress for creators. In several cases, these remarks appeared alongside strings of other taunts, suggesting that their impact is not just in their wording but in their cumulative pressure.
@@ V-D Humor, Satire, and Ambiguity in Model Moderation
One such comment was ”Can’t sing for nothing!!!!”, flagged by GPT despite being a common form of exaggerated critique in entertainment contexts. Similarly, ”2:33 too much jiggling” was flagged by Gemini, likely due to its focus on physical attributes, though the intent may have been commentary rather than harassment.
Comments referencing public figures or using layered sarcasm were also commonly misclassified. For instance, ”It sounds like UserNameProtected Cobain’s remembrance video” was flagged by GPT, Gemini, and Claude, though human reviewers judged it to be a stylistic comparison rather than an attack. Another flagged comment, ”You just gonna act like nothing happened?…” shows how models can mistake rhetorical or skeptical questions for aggression, especially when stripped of conversational context.
Claude and Gemini in particular flagged remarks like ”why does he look malnourished” or ”UserNameProtected ain’t even real” that are commonly used in meme culture or satire. These comments, while edgy, were not interpreted as harmful by human reviewers. However, the models appeared sensitive to tone and phrasing that hinted at ridicule or disbelief.
These cases underscore the limits of current models in interpreting intent, especially when humor blurs the line between critique and cruelty. Without access to conversational history or platform-specific norms, LLMs often err on the side of caution, flagging emotionally charged or socially playful content as harmful.
@@ VI Future Work
VI-A Thread-Level and Multi-modal Context
@@ VI-C Ensemble and Risk-Tiered Pipelines
VI-D User-Facing Feedback and Digital Citizenship
@@ VI-E Long-Term Outcomes and Mental-Health Impact
Finally, future studies should trace whether improved moderation actually lowers anxiety, self-harm ideation, or community churn. Partnering with mental-health researchers and using de-identified longitudinal data can turn model gains into real-world wellbeing.
@@ VII Conclusion
This paper benchmarks OpenAI GPT-4.1, Google Gemini 1.5 Pro, and Anthropic Claude 3 Opus on 5 080 YouTube comments sampled from high-abuse threads in gaming, lifestyle, food vlog, and music channels. Each model was evaluated with the same prompt under deterministic settings. Gemini identified the largest share of harmful content with recall 0.875 but its precision fell to 0.767 due to frequent false positives. Claude achieved the highest precision 0.920 and the lowest false-positive rate 0.022 yet its recall slipped to 0.720. GPT delivered the best overall balance with an F1 score 0.863, precision 0.887, and recall 0.841. Qualitative analysis showed that sarcasm, coded insults, and mixed-language slang remain persistent blind spots for all three models. These findings make clear that no single system meets every moderation requirement. Practical pipelines should combine complementary models, incorporate conversational context, and fine-tune for under-represented languages and subtle forms of abuse. The fully de-identified dataset and detailed prompts are publicly released to support reproducible research and foster further improvements in automated content moderation.

### v2

# Moderating Harm: Benchmarking Large Language Models for Cyberbullying Detection in YouTube Comments

@@ Abstract
As online platforms grow, comment sections increasingly host harassment that undermines user experience and well-being. This study benchmarks three leading large language models, OpenAI GPT 4.1, Google Gemini 1.5 Pro, and Anthropic Claude 3 Opus, on a corpus of 5 080 YouTube comments sampled from high abuse threads in gaming, lifestyle, food vlog, and music channels. The dataset comprises 1 334 harmful and 3 746 non-harmful messages in English, Arabic, and Indonesian, annotated independently by two reviewers with almost perfect agreement (Cohen’s $\kappa$ = 0.83). Using a unified prompt and deterministic settings, GPT-4.1 achieved the best overall balance with an F1 score of 0.863, precision of 0.887, and recall of 0.841. Gemini flagged the highest share of harmful posts (recall = 0.875) but its precision fell to 0.767 due to frequent false positives. Claude delivered the highest precision at 0.920 and the lowest false positive rate of 0.022, yet its recall dropped to 0.720. Qualitative analysis showed that all three models struggle with sarcasm, coded insults, and mixed-language slang. These results underscore the need for moderation pipelines that combine complementary models, incorporate conversational context, and fine-tune underrepresented languages and implicit abuse. A de-identified version of the dataset and full prompts is publicly released to promote reproducibility and further progress in automated content moderation.
@@
Artificial intelligence, cyberbullying, hate speech, large language models, natural language processing
@@ I Introduction
Online platforms now mediate much of everyday communication, with more than 4.95 billion active social media accounts worldwide in 2023 [1]. That ubiquity carries risk: recent studies report that between 6.5% and 35.4% of adolescents in the U.S. and Europe have experienced cyberbullying [2], and longitudinal studies link such exposure to elevated anxiety, depression, and suicidal ideation [3]. As the scale and psychological impact of online abuse grow, so does the urgency of developing reliable, context-aware moderation methods, especially for platforms popular with adolescents such as YouTube and TikTok.
@@ I-A Background
Cyberbullying has become a major digital safety concern because abusive material is persistent, anonymous, and can spread instantaneously across networks [4]. As global connectivity and social media adoption have expanded, online harassment has become more visible, with increasing concern about its scale and severity [5].
Comment sections on social networks, video sharing sites, and forums are particular hotspots. In a 2024 U.S. survey of parents, 79% said their children had encountered cyberbullying on YouTube [6]. A cross-national analysis of 180 000 adolescents in 42 countries also linked problematic social-media use to both cyberbullying victimization and perpetration, with stronger effects among girls [7].
@@ I-B Motivation and Problem Statement
This is especially dangerous in cyberbullying, where the impact of a single overlooked comment can be deeply personal [12]. Unlike hate speech or spam, bullying is often subtle: expressed through mockery, group pressure, or repeated jabs that may seem harmless in isolation but accumulate over time. Many academic benchmarks fail to capture this complexity, relying on synthetic or crowdsourced datasets with clear-cut abuse [13, 14].
This study addresses that gap by evaluating three most advanced LLMs, GPT-4, Gemini, and Claude, on authentic YouTube comments from videos with documented cyberbullying. By focusing on nuanced, emotionally loaded content, we examine not just whether models detect harmful language, but how well they handle ambiguity, cultural context, and tone.
@@ I-C Objectives of the Study
I-C1 Benchmark performance
Compare OpenAI GPT, Google Gemini, and Anthropic Claude in detecting cyberbullying within real YouTube data.
I-C2 Quantify error patterns
Measure false positive and false negative rates and identify recurring misclassification themes.
I-C3 Conduct qualitative error analysis
Manually inspect model outputs to assess contextual understanding and pinpoint linguistic edge cases.
I-C4 Assess robustness across nuance
Evaluate how well each model handles abusive content that is linguistically ambiguous, culturally specific, or indirectly expressed.
@@ II-A AI Content Moderation and Toxic Language Detection
However, most benchmarks are composed of isolated sentences and short comments labeled through crowdwork. They rarely reflect the conversational, multi-turn, or emotionally charged dynamics seen in real-world platforms. Furthermore, these datasets often over-represent overt toxicity while under-representing gray area content such as mockery, dismissive sarcasm, or personal digs. As a result, models trained on them perform well in benchmark settings but struggle in live deployments where language is more ambiguous and emotionally expressive [18, 19].
@@ II-B Challenges in Detecting Implicit and Nuanced Abuse
Detecting implicit harm remains one of the most difficult challenges in AI-based moderation. Sarcasm, coded language, cultural slang, and subtextual insults frequently evade model detection, even in LLMs. For example, a phrase like “She’s always such a queen, right?” may be an insult depending on tone and context, but appears harmless in isolation. Studies in sarcasm detection [20], adversarial misspellings, irony detection, and tone modeling has attempted to address this gap, but progress remains limited [21].
Recent research also suggests that models often over-flag emotionally intense or sensitive language even when used in supportive contexts, leading to user frustration and moderation fatigue. This tension between under-flagging subtle harm and over-flagging emotional expression makes fine-tuned moderation especially difficult. Studies have proposed integrating tone analysis, conversational history [22], and speaker intent as ways to improve subtle abuse detection, but these are rarely deployed at scale.
II-C Moderating Multilingual and Under-resourced Content
While English dominates most training corpora, abuse occurs in every language, and much of it is multilingual, transliterated, or mixed with emojis and slang. Studies consistently show significant drops in model accuracy when applied to languages such as Arabic [23] and Hindi [24], particularly when users employ spelling variation or obfuscation to evade filters [25]. Code-switching and transliteration further degrade detection performance, as models frequently miss common abusive patterns or misinterpret them out of context [26, 27].
Cross-lingual models and prompt-tuned systems have demonstrated moderate performance improvements, especially when fine-tuned on translated or augmented datasets [28]. However, significant language imbalance remains, particularly in online spaces with small yet highly active non-English user communities [29]. For platforms with global audiences, this poses a critical equity issue: harmful content in English is far more likely to be effectively moderated than equivalent abuse in low-resource languages or regional dialects [30].
@@ II-D Evaluating LLMs for Safety and Fairness in Moderation
The rise of general-purpose LLMs such as GPT-4, Claude, and Gemini has driven increasing interest in their application to content moderation tasks. Recent research indicates that these models can outperform traditional task-specific classifiers in zero-shot or few-shot scenarios, especially on previously unseen or nuanced content [31]. Their ability to dynamically incorporate platform policies, contextual reasoning, or nuanced moderation guidelines directly into their prompting strategies enhances their flexibility and adaptability [32].
However, the performance of LLMs in moderation tasks is highly sensitive to prompt phrasing, target domains, and cultural nuances [33, 34]. Sociocultural audit frameworks have recently been proposed, employing persona-based prompts [35] and synthetic demographic simulations to systematically evaluate model fairness and reveal hidden biases or blind spots [36, 37]. These frameworks highlight how LLM responses may shift significantly based on perceived user identity or topical framing, raising substantial concerns around fairness, consistency, and equitable moderation [38].
While LLMs offer substantial advantages in scalability, adaptability, and reduced reliance on explicit rule-based configurations [39], their inherent dependence on pretraining data, which often contain historical biases, and the opacity of their decision-making processes continue to present barriers to safe and transparent deployment in moderation workflows [40].
@@ III-B Data Cleaning and Anonymization
Each comment underwent a two–step preprocessing pipeline. First, text was normalized (UTF-8 decoding, whitespace trimming, emoji and punctuation were preserved). Second, personally identifiable information including user names, real names, phone numbers, e-mail addresses, links, and explicit geo-markers was either deleted or replaced by neutral placeholders such as “UserNameProtected”. No other linguistic content was altered. This procedure leaves the semantic core of each message intact while mitigating re-identification risk, allowing the corpus to be shared for research without exposing private data.
@@ III-C2 Not harmful
Inter‐rater reliability was high, with 91.0% raw agreement and $\kappa=0.83$, which is considered almost perfect. [^1] Landis and Koch, 1977. Disagreements (9% of instances) were adjudicated by discussion until consensus, and that label was written to the final ground truth file. Labels were never exposed to the models.
@@ III-G Manual Review and Model Comparison
Quantitative scores alone do not reveal why a system succeeds or fails, so we carried out a post-hoc qualitative review. Two annotators independently inspected a stratified sample of 200 disagreements, consisting of 50 false positives, 50 false negatives, and 100 edge case ties, drawn in equal proportion from all three models. For each comment we discussed (i) whether the human ground truth should stand and (ii) what linguistic cues might have misled the model (sarcasm, coded slurs, emoji, topic drift, and so forth). Notes from this exercise were grouped into recurring error themes that inform the Discussion section. The procedure does not alter any numeric results but clarifies how each system handles nuance, context, and cultural references.
@@ IV-A Dataset Composition
The evaluation dataset consisted of 5 080 comments, including 1 334 harmful and 3 746 non-harmful instances. Harmful content made up about 26.3% of the total dataset, indicating a moderate class imbalance that could influence model metrics like precision and recall. Table I provides the class breakdown.
@@ V-A False Positives and Surface-Level Flagging
To better understand model over-flagging, we analyzed cases where AI models marked comments as harmful while human reviewers did not. These false positives were grouped by model agreement patterns to highlight shared tendencies and specific weaknesses. While most flagged content reflected strong language or emotional tone, it often lacked any abusive intent. Below are four common patterns observed.
V-A1 All Three Models vs. Human: Shared Over-flagging Patterns
@@ V-A3 Gemini: Over-Flagging Concern and Instability
Gemini frequently flagged comments that referenced mental health, personal change, or emotional concern, even when human reviewers found no clear harm. Many of these comments offered support, expressed worry, or questioned the authenticity of a user’s behavior in dramatic or sarcastic terms. Phrases like “please take care of yourself” or “you can’t keep scaring everyone like this” were often interpreted by Gemini as harmful. This suggests a tendency to over-flag emotionally sensitive or speculative language, particularly when it touches on perceived instability or trauma.
@@ V-C Risks of Missed Moderation in Sensitive Cases
More troubling were coordinated suggestions to “unfollow so he falls off” or “ignore him until he cracks again.” These comments reflect collective hostility rather than isolated aggression. When such messages go unflagged, they enable mob harassment campaigns that can lead to reputational harm, social withdrawal, or emotional distress for creators. In several cases, these remarks appeared alongside strings of other taunts, suggesting that their impact is not just in their wording but in their cumulative pressure.
@@ V-D Humor, Satire, and Ambiguity in Model Moderation
One such comment was “Can’t sing for nothing!!!!”, flagged by GPT despite being a common form of exaggerated critique in entertainment contexts. Similarly, “2:33 too much jiggling” was flagged by Gemini, likely due to its focus on physical attributes, though the intent may have been commentary rather than harassment.
Comments referencing public figures or using layered sarcasm were also commonly misclassified. For instance, “It sounds like UserNameProtected Cobain’s remembrance video” was flagged by GPT, Gemini, and Claude, though human reviewers judged it to be a stylistic comparison rather than an attack. Another flagged comment, “You just gonna act like nothing happened?…” shows how models can mistake rhetorical or skeptical questions for aggression, especially when stripped of conversational context.
Claude and Gemini in particular flagged remarks like “UserNameProtected ain’t even real” that are commonly used in meme culture or satire. These comments, while edgy, were not interpreted as harmful by human reviewers. However, the models appeared sensitive to tone and phrasing that hinted at ridicule or disbelief.
These cases underscore the limits of current models in interpreting intent, especially when humor blurs the line between critique and cruelty. Without access to conversational history or platform-specific norms, LLMs often make conservative judgments, flagging emotionally charged or socially playful content as harmful.
@@ VI Future Work
VI-A Thread-Level and Multimodal Context
@@ VI-C Ensemble and Risk-Tiered Pipelines
VI-D User-Facing Feedback
@@ VI-E Long-Term Outcomes and Mental-Health Impact
Finally, future studies should trace whether improved moderation actually lowers anxiety, self-harm ideation, or community churn. Partnering with mental-health researchers and using de-identified longitudinal data can turn model gains into real-world well-being.
@@ VII Conclusion
This paper benchmarks OpenAI GPT-4.1, Google Gemini 1.5 Pro, and Anthropic Claude 3 Opus on 5 080 YouTube comments sampled from high-abuse threads in gaming, lifestyle, food vlog, and music channels. Each model was evaluated with the same prompt under deterministic settings. Gemini identified the largest share of harmful content with recall 0.875 but its precision fell to 0.767 due to frequent false positives. Claude achieved the highest precision 0.920 and the lowest false positive rate 0.022 yet its recall slipped to 0.720. GPT delivered the best overall balance with an F1 score 0.863, precision 0.887, and recall 0.841. Qualitative analysis showed that sarcasm, coded insults, and mixed-language slang remain persistent blind spots for all three models. These findings make clear that no single system meets every moderation requirement. Practical pipelines should combine complementary models, incorporate conversational context, and fine-tune for under-represented languages and subtle forms of abuse. The fully de-identified dataset and detailed prompts are publicly released to support reproducible research and foster further improvements in automated content moderation.